Show that 2 power 0 is equal to 1


1
Expert's answer
2021-02-10T10:09:16-0500

2 power 0 is equal to 1 by definition, if we want to use power rules. For example, let's divide "2^n" by "2^n" where "n" is any number. It is clear, that


"\dfrac{2^n}{2^n} = 1"

But, on the other hand, using power rules, obtain:


"\dfrac{2^n}{2^n} = 2^{n-n} = 2^0"

Thus for the sake of consistency it should be "2^0 = 1".